OpenStudy (anonymous):

Ok so this question just really confuses me. I know you just do chain rule a bunch but I am still lost. Question below

OpenStudy (anonymous):

\[\sqrt{x+\sqrt{x+\sqrt{x}}}\] Find the derivative.

OpenStudy (noelgreco):

First, eliminate the radicals - use fractions and grouping symbols.